# Dynamic Contextual Pricing with Doubly Non-Parametric Random Utility Models

### v1

@@ 2.1 Mean Utility Function Estimation
To address the challenge of not directly observing random utility $v_{t}$, we observe the following property for both single-product dynamic pricing. Specifically, if we adopt the uniform random policy $p_{t}\sim\operatorname{Unif}(0,B)$, we can recover the mean utility function in an average sense. Denote by $g_{t}:=By_{t}$. It can be observed that given the contextual covariate $\boldsymbol{x}_{t}$, $g_{t}$ satisfies
$\mathbb{E}[g_{t}\,|\,\boldsymbol{x}_{t}]$ $=\mathbb{E}_{\varepsilon_{t}}\left[\mathbb{E}_{p_{t}}[By_{t}\,|\,\boldsymbol{x}_{t},\varepsilon_{t}]\;|\;\boldsymbol{x}_{t}\right]$ (5)
@@
To non-parametrically estimate the mean utility function $\mathring{\mu}(\cdot)$, we apply the method of Distributional Nearest Neighbors (DNN, Steele (2009); Biau et al. (2010)) which is flexible and applicable even with a limited amount of data. DNN is a variation of $K$-Nearest Neighbours and automatically assigns monotonic weights to the nearest neighbors in a distributional fashion on the entire sample. It incorporates the idea of bagging by averaging all 1-NN estimators, each constructed from randomly subsampling $s$ observations without replacement. Here $s$ is the subsampling scale diverging with the total sample size.
Suppose a sample $\boldsymbol{z}_{t}=\{(\boldsymbol{x}_{t},g_{t})\}_{t\in{\cal T}_{k,exp}}$ of size $n_{k,exp}$ is collected in the exploration phase of $k$th episode. Considering a new contextual covariate observation $\boldsymbol{x}$, we would like to fit a non-parametric mean function from the samples to estimate the underlying truth $\mathring{\mu}(\boldsymbol{x})$. For $1\leq s\leq n_{k,exp}$, the DNN estimator $\widehat{\mu}_{k}^{\operatorname{DNN}}(\boldsymbol{x};s)$ with sub-sampling scale $s$ for estimating $\mathring{\mu}(\boldsymbol{x})$ is formally defined as the following $U$-statistic
@@
where $g_{(1)}(\boldsymbol{z}_{i_{1}},\cdots,\boldsymbol{z}_{i_{s}})$ is the $1$-NN estimator for estimating $\mathring{\mu}(\boldsymbol{x})$ based on the subsample $(\boldsymbol{z}_{i_{1}},\cdots,\boldsymbol{z}_{i_{s}})$. Formally, it returns the value of $g_{(1)}(\boldsymbol{z}_{i_{1}},\cdots,\boldsymbol{z}_{i_{s}})=g_{(1)}$ from the reordered sample $\{(\boldsymbol{x}_{(1)},g_{(1)}),\cdots,(\boldsymbol{x}_{(s)},g_{(s)})\}$ of $\{(\boldsymbol{x}_{i_{1}},g_{i_{1}}),\ldots,(\boldsymbol{x}_{i_{s}},g_{i_{s}})\}$ such that
$\left\lVert\boldsymbol{x}_{(1)}-\boldsymbol{x}\right\rVert\leq\left\lVert\boldsymbol{x}_{(2)}-\boldsymbol{x}\right\rVert\leq\cdots\leq\left\lVert\boldsymbol{x}_{(s)}-\boldsymbol{x}\right\rVert,$ (7)
[…]
@@
where $\{(\boldsymbol{x}_{(1)},g_{(1)}),\ldots,(\boldsymbol{x}_{(n_{k,exp})},g_{(n_{k,exp})})\}$ is the overall ordered sample such that
$\left\lVert\boldsymbol{x}_{(1)}-\boldsymbol{x}\right\rVert\leq\left\lVert\boldsymbol{x}_{(2)}-\boldsymbol{x}\right\rVert\leq\cdots\leq\left\lVert\boldsymbol{x}_{(n_{k,exp})}-\boldsymbol{x}\right\rVert.$
A nice feature of the above DNN estimator is that the weight distribution is characterized exclusively by the total sample size $n_{k,exp}$ and the subsampling scale $s$. Therefore, the L-statistic representation in (8) can be exploited for easy implementation, while the U-statistic representation is useful in theoretical analysis for the uniform convergence rate of the estimator.
@@
When the mean utility function and density function of covariates $\boldsymbol{x}_{t}$ have a higher level of smoothness, two-scale DNN (TDNN, Demirkaya et al. (2022)) can be applied to remove the first-order bias of the DNN estimator by aggregating two DNN estimators with distinct subsampling scales $s_{1}$ and $s_{2}$. Specifically, a TDNN estimator of the mean utility function $\mathring{\mu}(\boldsymbol{x})$ is given by
@@
In general, we can propose an $m$-scale DNN estimator to eliminate the bias up to the $(m-1)$-th order by constructing a linear combination of $m$ DNN estimators with distinct subsampling scales, utilizing a similar technique in (9).
@@ 2.2 Utility Noise Distribution Estimation
Thus, $F(\cdot)$ can be estimated by the Nadaraya-Watson kernel regression estimator. Given sample $\{\boldsymbol{x}_{t},y_{t},p_{t}\}_{t\in{\cal T}_{k,exp}}$ collected in the exploration phase and any estimated utility function $\mu(\cdot)$, we define
@@
$\widehat{F}_{k}\left(z;\widehat{\mu}_{k}^{\operatorname{DNN}}\right)==1-\frac{\widehat{a}_{k}(z;\widehat{\mu}_{k}^{\operatorname{DNN}})}{\widehat{\xi}_{k}(z;\widehat{\mu}_{k}^{DNN})}.$ (13)
Similarly, the Nadaraya-Watson estimator of $F(z)$ based on the TDNN estimator of the mean utility function is given by
$\widehat{F}_{k}\left(z;\widehat{\mu}_{k}^{\operatorname{TDNN}}\right)==1-\frac{\widehat{a}_{k}(z;\widehat{\mu}_{k}^{\operatorname{TDNN}})}{\widehat{\xi}_{k}(z;\widehat{\mu}_{k}^{\operatorname{TDNN}})}.$ (14)
@@ 2.3 Optimal Decision
$p_{t}^{*}=h\circ\mathring{\mu}(\boldsymbol{x}_{t}),$ (16)
$\text{where}\quad h(v):=v+\phi^{-1}(-v)\quad\mbox{and}\quad\phi(z):=z-\frac{1-F(z)}{F^{(1)}(z)}.$
Here $h\circ\mathring{\mu}(\cdot)$ denotes the composite function $h(\mathring{\mu}(\cdot))$ and $F^{(1)}$ is the first-order derivative of $F$.
@@
In reality, we can only estimate $\mathring{\mu}(\cdot)$ and $F(\cdot)$ from collected data. In the $k$-th episode of Algorithm 1, we obtain the estimated mean utility function $\widehat{\mu}_{k}$ using either the DNN or TDNN estimator and calculate the Nadaraya-Watson estimator $\widehat{F}_{k}(z)$ based on (13) or (14). Accordingly, the data-driven optimal offer price $\widehat{p}_{t}$ for $t\in{\cal T}_{k,com}$ in the exploitation phase is obtained by optimizing the estimated reward. That is, $\widehat{p}_{t}=\underset{p_{t}>0}{\operatorname{argmax}}~{}p_{t}[1-\widehat{F}_{k}(p_{t}-\widehat{\mu}_{k}(\boldsymbol{x}_{t})]$. As a result, we have that for any $t\in{\cal T}_{k,com}$,
@@ 3 Theoretical Guarantees
In this section, we establish the regret analysis of the proposed Algorithm 1 for doubly robust learning and decision in contextual dynamic pricing. For the theoretical analysis, we first introduce some notations and present necessary assumptions. For any candidate mean utility function $\mu:\mathcal{X}\mapsto\mathbb{R}$, let $\xi(z;\mu)$ be the density function of $Z=p_{t}-\mu(\boldsymbol{x}_{t})$ evaluated at $Z=z$ and define $F(z;{\mu}):=1-\mathbb{E}[y_{t}|p_{t}-{\mu}(\boldsymbol{x}_{t})=z]$. Therefore, the underlying true noise distribution $F(z)=F(z;\mathring{\mu})$ according to (10). We next introduce the assumptions about the market noise $\varepsilon_{t}$, the distribution function $F(u;\mu)$, density function $\xi(u;\mu)$, and the kernel function $K(x)$.
@@
There exists positive constants $L_{\phi}$ and $B_{u}$ such that $\phi^{(1)}(z)\geq L_{\phi}$ for all $z\in{\cal S}_{\varepsilon}$ with $\phi(z)$ defined in (16).
Assumptions 1, 2, and 3 are standard assumptions for the smoothness of the relevant functions in nonparametric estimation and inference. The rate of convergence $(T^{-\frac{2m+1}{(4m-1)(d+2)}}\land T^{-\frac{1}{2(d+4)}})\sqrt{d\log T}$ in Assumption 2 can be satisfied for the DNN or TDNN estimators (with a further improved neighborhood size) with appropriately chosen parameters. Assumption 4 guarantees that $\phi(x)$ is strictly increasing and, thus, the uniqueness of the optimal solution of (15). It is weaker than that in Javanmard and Nazerzadeh (2019) since $1-F(u)$ is log-concave under the further restriction of $L_{\phi^{\prime}}\geq 1$.
The regret analysis unfolds in two steps. Firstly, we will develop the uniform convergence rate of both the DNN and TDNN estimator in estimating the mean utility function $\mathring{\mu}$, as well as the convergence rate of the Nadaraya-Watson estimator for estimating the noise distribution $F(\cdot)$. Secondly, building upon the uniform convergence rate of the estimators, we will obtain an upper bound for the regret of Algorithm 1.
@@ 3.1 Statistical Convergence Rate of Learning
Suppose density function $f(\cdot)$ of $\boldsymbol{x}_{t}$ and the mean utility function $\mathring{\mu}(\cdot)$ are two times continuously differentiable with bounded second-order derivatives on the compact support $\mathcal{X}$ of $\boldsymbol{x}_{t}$. The density $f(\boldsymbol{x})$ is bounded away from 0 and $\infty$ for $\boldsymbol{x}\in\mathcal{X}$. Then we have for the DNN estimator that with probability $1-2\delta$,
@@
The uniform convergence rate results in theorem 5, and its proof is new to the literature and is of independent interest. Regarding the the uniform convergence of DNN estimator, the first error term $s^{-2/d}$ provides an upper bound for the bias, while the second error term corresponds to the variance. In a similar manner, in the uniform bound for TDNN estimator, the first error term $s_{2}^{-4/d}$ corresponds to the bias and the second error term is associated with the variance. We can observe from Theorem 5 that the TDNN estimator has the same order of variance as the DNN estimator but reduces the bias. This improvement is achieved by TDNN’s formulation of a linear combination of two DNN estimators aimed at eliminating the first-order bias. For any general integer $m\geq 2$, one can employ an $m$-scale DNN estimator, which is defined as a linear combination of $m$ DNN estimators, each with a distinct subsampling scale to remove the bias up to the $(m-1)$-th order. This requires a higher-order of smoothness for the mean utility function and density function of $\boldsymbol{x}_{t}$. Moreover, it brings the complexity of tuning $m$ distinct subsampling scales and the risk of increased variance.
@@
Assume DNN estimator $\widehat{\mu}_{k}^{\operatorname{DNN}}$ is applied with subsampling size $s=O(n_{k,\exp}^{\frac{d}{d+4}})$. Under condition of Theorem 5 for DNN estimator and Assumptions 14, there exist constants $C_{1}>0$ and $C_{2}>0$ such that for any episode $k$ satisfying $n_{k,exp}^{\frac{2m}{2m+1}}\geq\max(C_{1}d(\frac{\log n_{k,exp}}{2m+1}+1),3\log n_{k,exp})$ with probability at least $1-2\delta$ with $\delta\in(4\exp\{-C_{2}n_{k,exp}^{\frac{m}{2m+1}}/\log n_{k,exp}\},1/2]$, it holds that
@@
The first term in the upper bound comes from kernel approximation of the noise distribution, while the second term comes from the upstream estimation error of DNN estimator for mean utility $\mathring{\mu}(\boldsymbol{x})$. For existing work that assumes a parametric model of $\mathring{\mu}(\boldsymbol{x})$, the second error term is dominated by the first term (Fan et al. (2022)). However, Theorem 6 shows that the upstream estimation error of $\mathring{\mu}(\boldsymbol{x})$ becomes non-negligible as the dimensionality $d$ grows under the more general non-parametric setting. A parallel result is also established in Appendix LABEL:TDNN-parallel-result of the supplementary material for the kernel estimator $\widehat{F}_{k}\left(z;\widehat{\mu}_{k}^{\operatorname{TDNN}}\right)$ when TDNN estimator is applied to estimate the mean utility function.

### v2

@@ 2.1 Mean Utility Function Estimation
To address the challenge of not directly observing random utility $v_{t}$, we observe the following property for both single-product dynamic pricing. Specifically, if a uniform random policy $p_{t}\sim\operatorname{Unif}(0,B)$ with $B$ being the upper bound for the market value is adopted, we can recover the mean utility function in an average sense. Denote by $g_{t}:=By_{t}$. It can be observed that given the contextual covariate $\boldsymbol{x}_{t}$, $g_{t}$ satisfies
$\mathbb{E}[g_{t}\,|\,\boldsymbol{x}_{t}]$ $=\mathbb{E}[By_{t}\,|\,\boldsymbol{x}_{t}]=\mathbb{E}_{\varepsilon_{t}}\left[\mathbb{E}_{p_{t}}[By_{t}\,|\,\boldsymbol{x}_{t},\varepsilon_{t}]\;|\;\boldsymbol{x}_{t}\right]$ (5)
@@
To non-parametrically estimate the mean utility function $\mathring{\mu}(\cdot)$, we apply the method of Distributional Nearest Neighbors (DNN, Steele (2009); Biau et al. (2010)) which is flexible and applicable even with a limited amount of data. DNN is a variation of $K$-Nearest Neighbours ($K$-NN) and automatically assigns monotonic weights to the nearest neighbors in a distributional fashion on the entire sample. It incorporates the idea of bagging by averaging all 1-NN estimators, each constructed from randomly subsampling $s$ observations without replacement. Here $s$ is the subsampling scale diverging with the total sample size.
Suppose a sample $\{(\boldsymbol{x}_{t},y_{t})\}_{t\in{\cal T}_{k,exp}}$ of size $n_{k,exp}$ is collected in the exploration phase of the $k$th episode. Since $g_{t}:=By_{t}$ satisfies (5), we build our estimator directly with $\{(\boldsymbol{x}_{t},g_{t})\}_{t\in{\cal T}_{k,exp}}$. Considering a new contextual covariate observation $\boldsymbol{x}$, our goal is to fit a non-parametric mean function from the sample to estimate the underlying truth $\mathring{\mu}(\boldsymbol{x})$. We reorder the observations $\{(\boldsymbol{x}_{t},g_{t})\}_{t\in{\cal T}_{k,exp}}$ according to $\boldsymbol{x}_{t}$’s Euclidean distance to $\boldsymbol{x}$ and denote the reordered sample as $\{(\boldsymbol{x}_{(1)},g_{(1)}),\ldots,(\boldsymbol{x}_{(n_{k,exp})},g_{(n_{k,exp})})\}$ such that
$\|\boldsymbol{x}_{(1)}-\boldsymbol{x}\|\leq\|\boldsymbol{x}_{(2)}-\boldsymbol{x}\|\leq\cdots\leq\|\boldsymbol{x}_{(n_{k,exp})}-\boldsymbol{x}\|,$ (6)
[…]
@@
where $g_{(1)}(\boldsymbol{z}_{i_{1}},\cdots,\boldsymbol{z}_{i_{s}})$ is the $1$-NN estimator of $\mathring{\mu}(\boldsymbol{x})$ based on the subsample $(\boldsymbol{z}_{i_{1}},\cdots,\boldsymbol{z}_{i_{s}})$. Specifically, with a abuse of notation, $g_{(1)}(\boldsymbol{z}_{i_{1}},\cdots,\boldsymbol{z}_{i_{s}}):=g_{(1)}$ from the reordered sample $\{(\boldsymbol{x}_{(1)},g_{(1)}),\cdots,(\boldsymbol{x}_{(s)},g_{(s)})\}$ of the subsample $\{(\boldsymbol{x}_{i_{1}},g_{i_{1}}),\ldots,(\boldsymbol{x}_{i_{s}},g_{i_{s}})\}$ such that
$\|\boldsymbol{x}_{(1)}-\boldsymbol{x}\|\leq\|\boldsymbol{x}_{(2)}-\boldsymbol{x}\|\leq\cdots\leq\|\boldsymbol{x}_{(s)}-\boldsymbol{x}\|.$ (8)
The subsampling scale $s$ will be chosen to balance the trade-off between bias and variance. In addition, it is shown in Steele (2009) that the DNN estimator also admits an equivalent representation of a weighted nearest neighbor that
@@
where $\{(\boldsymbol{x}_{(1)},g_{(1)}),\ldots,(\boldsymbol{x}_{(n_{k,exp})},g_{(n_{k,exp})})\}$ is the overall ordered sample according to the ordered distance in (6).
A nice feature of the above DNN estimator is that the weights on the nearest neighbors are characterized exclusively by the total sample size $n_{k,exp}$ and the subsampling scale $s$. Therefore, the equivalent representation in (9) can be exploited for easy implementation, while the representation in (7) is crucial in theoretical analysis for the uniform convergence rate of the estimator.
@@
When the mean utility function and density function of covariates $\boldsymbol{x}_{t}$ have a higher degree of smoothness, two-scale DNN (TDNN, Demirkaya et al. (2022)) can be applied to eliminate the first-order bias of the DNN estimator by aggregating two DNN estimators with distinct subsampling scales $s_{1}$ and $s_{2}$. Specifically, a TDNN estimator of the mean utility function $\mathring{\mu}(\boldsymbol{x})$ is given by
@@
In general, if the mean utility function $\mathring{\mu}(\boldsymbol{x})$ and density function of covariates $\boldsymbol{x}_{t}$ are ($2l$)-times differentiable for an integer $l\geq 2$, we can construct an $l$-scale DNN estimator to eliminate the bias up to the $(l-1)$-th order by constructing a linear combination of $l$ DNN estimators with distinct subsampling scales, utilizing a similar technique in (10).
@@ 2.2 Utility Noise Distribution Estimation
Thus, the Nadaraya-Watson kernel regression estimator can be applied to estimate $F(\cdot)$. Given sample $\{\boldsymbol{x}_{t},y_{t},p_{t}\}_{t\in{\cal T}_{k,exp}}$ collected in the exploration phase and any estimated utility function $\mu(\cdot)$, we define
@@
$\widehat{F}_{k}\left(z;\widehat{\mu}_{k}^{\operatorname{DNN}}\right)=1-\frac{\widehat{a}_{k}(z;\widehat{\mu}_{k}^{\operatorname{DNN}})}{\widehat{\xi}_{k}(z;\widehat{\mu}_{k}^{DNN})}.$ (14)
Analogously, the Nadaraya-Watson estimator of $F(z)$ based on the TDNN estimator of the mean utility function is given by
$\widehat{F}_{k}\left(z;\widehat{\mu}_{k}^{\operatorname{TDNN}}\right)=1-\frac{\widehat{a}_{k}(z;\widehat{\mu}_{k}^{\operatorname{TDNN}})}{\widehat{\xi}_{k}(z;\widehat{\mu}_{k}^{\operatorname{TDNN}})}.$ (15)
@@ 2.3 Optimal Decision
$p_{t}^{*}=h\circ\mathring{\mu}(\boldsymbol{x}_{t}),\text{where}\quad h(v)$ $:=v+\phi^{-1}(-v)\quad\mbox{and}\quad\phi(z):=z-\frac{1-F(z)}{F^{(1)}(z)}.$ (17)
Here $h\circ\mathring{\mu}(\cdot)$ is the composite function $h(\mathring{\mu}(\cdot))$ and $F^{(1)}$ is the first-order derivative of $F$.
@@
In reality, we can only estimate $\mathring{\mu}(\cdot)$ and $F(\cdot)$ from collected data. In the $k$-th episode of Algorithm 1, we obtain the estimated mean utility function $\widehat{\mu}_{k}$ using either the DNN or TDNN estimator and calculate the kernel estimator $\widehat{F}_{k}(z)$ based on (14) or (15). Accordingly, the data-driven optimal offered price $\widehat{p}_{t}$ for $t\in{\cal T}_{k,com}$ in the exploitation phase is obtained by optimizing the estimated reward. That is, $\widehat{p}_{t}=\underset{p_{t}>0}{\operatorname{argmax}}~{}p_{t}[1-\widehat{F}_{k}(p_{t}-\widehat{\mu}_{k}(\boldsymbol{x}_{t})]$. As a result, it can be obtained that for any $t\in{\cal T}_{k,com}$,
@@ 3 Theoretical Guarantees
In this section, we establish the regret analysis of the proposed Algorithm 1 for doubly robust learning and decision in contextual dynamic pricing. For the theoretical analysis, we first introduce some notations and present necessary assumptions. For any candidate mean utility function $\mu:\mathcal{X}\mapsto\mathbb{R}$, let $\xi(z;\mu)$ be the density function of $Z=p_{t}-\mu(\boldsymbol{x}_{t})$ evaluated at $Z=z$ and define $F(z;{\mu}):=1-\mathbb{E}[y_{t}|p_{t}-{\mu}(\boldsymbol{x}_{t})=z]$. Therefore, the underlying true noise distribution $F(z)=F(z;\mathring{\mu})$ according to (11). We next introduce the assumptions about the market noise $\varepsilon_{t}$, the distribution function $F(z;\mu)$, density function $\xi(z;\mu)$, and the kernel function $K(x)$.
@@
There exist positive constants $L_{\phi}$ and $B_{u}$ such that $\phi^{(1)}(z)\geq L_{\phi}$ for all $z\in{\cal S}_{\varepsilon}$ with $\phi(z)$ defined in (17).
Assumptions 1, 2, and 3 are standard assumptions for the smoothness of the relevant functions in nonparametric estimation and inference. The rate of convergence $(T^{-\frac{2m+1}{(4m-1)(d+2)}}\land T^{-\frac{1}{2(d+4)}})\sqrt{d\log T}$ in Assumption 2 can be satisfied for the DNN estimator or TDNN estimator (with a further shrinked neighborhood size) with appropriately chosen subsampling sizes. Assumption 4 guarantees that $\phi(x)$ is strictly increasing and, thus, the uniqueness of the optimal solution of (16). It is weaker than that in Javanmard and Nazerzadeh (2019) since $1-F(u)$ is log-concave under the further restriction of $L_{\phi^{\prime}}\geq 1$.
The regret analysis unfolds in two steps. First, we will develop the uniform convergence rate of both the DNN and TDNN estimators in estimating the mean utility function $\mathring{\mu}$, as well as the convergence rate of the Nadaraya-Watson kernel estimator for estimating the noise distribution $F(\cdot)$. Second, building upon the uniform convergence rates established in the first step, we will derive an upper bound for the regret of Algorithm 1.
@@ 3.1 Statistical Convergence Rate of Learning
Suppose density function $f(\cdot)$ of $\boldsymbol{x}_{t}$ and the mean utility function $\mathring{\mu}(\cdot)$ are twice continuously differentiable with bounded second derivatives on the compact support $\mathcal{X}$ of $\boldsymbol{x}_{t}$. The density $f(\boldsymbol{x})$ is bounded away from 0 and $\infty$ for $\boldsymbol{x}\in\mathcal{X}$. Then we have for the DNN estimator that with probability $1-2\delta$,
@@
The proof of Theorem 5 can be found in Section C of the supplementary material. The uniform convergence rate results in theorem 5 and the proofs are new to the literature and hold independent interest. Regarding the the uniform convergence of DNN estimator, the first error term $s^{-2/d}$ provides an upper bound for the bias, while the second error term corresponds to bound for the variance. In a similar manner, in the uniform bound for TDNN estimator, the first error term $s_{2}^{-4/d}$ is related to the bias and the second error term is associated with the variance. We can observe from Theorem 5 that the TDNN estimator preserves the same order of variance as the DNN estimator but reduces the bias. This improvement is achieved by TDNN’s formulation of a linear combination of two DNN estimators, targeted at eliminating the first-order bias. For any general integer $l\geq 2$, one can employ an $l$-scale DNN estimator, which is defined as a linear combination of $l$ DNN estimators, each with a distinct subsampling scale to remove the bias up to the $(l-1)$-th order. This construction requires the mean utility function and density function of $\boldsymbol{x}_{t}$ to have bounded derivatives up to the $2l$-th order.
@@
Assume DNN estimator $\widehat{\mu}_{k}^{\operatorname{DNN}}$ is applied with subsampling size $s=O(n_{k,\exp}^{\frac{d}{d+4}})$. Under condition of Theorem 5 for DNN estimator and Assumptions 14, there exist constants $C_{1}>0$ and $C_{2}>0$ such that for any episode $k$ satisfying $n_{k,exp}^{\frac{2m}{2m+1}}\geq\max(C_{1}d(\frac{\log n_{k,exp}}{2m+1}+1),3\log n_{k,exp})$ and $\delta\in(4\exp\{-C_{2}n_{k,exp}^{\frac{m}{2m+1}}/\log n_{k,exp}\},\frac{1}{2}]$, it holds with probability at least $1-2\delta$ that
@@
Theorem 6 follows directly from applying Lemma 11 in Appendix A and Theorem 5. The first term in the upper bound comes from kernel approximation of the noise distribution, while the second term comes from the upstream estimation error of DNN estimator for mean utility $\mathring{\mu}(\boldsymbol{x})$. For existing work that assumes a parametric model of $\mathring{\mu}(\boldsymbol{x})$, the second error term is dominated by the first term (Fan et al. (2022)). However, Theorem 6 shows that the upstream estimation error of $\mathring{\mu}(\boldsymbol{x})$ becomes non-negligible as the dimensionality $d$ grows under the more general non-parametric setting.
When TDNN estimator is applied to estimate the nonparametric mean regression function $\mathring{\mu}(\cdot)$ and then the same Nadaraya-Watson kernel estimator for the error distribution $F(\cdot)$, we can obtain the following parallel result in Theorem 7 below for the kernel estimator $\widehat{F}_{k}\left(z;\widehat{\mu}_{k}^{\operatorname{TDNN}}\right)$ when TDNN estimator is applied to estimate the mean utility function.
Theorem 7.
Assume TDNN estimator $\widehat{\mu}_{k}^{\operatorname{TDNN}}$ is applied with subsampling scales $s_{1}=O(n_{k,\exp}^{\frac{d}{d+8}\lor\frac{1}{7}})$ and $s_{2}\asymp s_{1}$. Under condition of Theorem 5 for TDNN estimator and Assumptions 14, there exist constants $C_{1}>0$ and $C_{2}>0$ such that for any episode $k$ satisfying $n_{k,exp}^{\frac{2m}{2m+1}}\geq\max(C_{1}d(\frac{\log n_{k,exp}}{2m+1}+1),3\log n_{k,exp})$ and $\delta\in(4\exp\{-C_{2}n_{k,exp}^{\frac{m}{2m+1}}/\log n_{k,exp}\},\frac{1}{2}]$, it holds with probability at least $1-2\delta$ that
$\sup_{z\in{\cal S}_{\varepsilon},{\mu}\in\mathcal{N}_{k}}\left\lvert\widehat{F}_{k}\left(z;\widehat{\mu}_{k}^{\operatorname{TDNN}}\right)-F(z)\right\rvert$ $\leq B_{F}n_{k,exp}^{-\frac{m}{2m+1}}\sqrt{\log n_{k,exp}}(\sqrt{d}+\sqrt{\log\delta^{-1}})$ (24)
$+CB_{F}{n_{k,exp}^{-(\frac{4}{d+8}\land\frac{3}{7})}\sqrt{d\log n_{k,exp}}}.$